OpenStudy (anonymous):

find the linearization L(x) of the function at a f(x) = lnx, a=1

OpenStudy (anonymous):

General form of linearization: y=f '(a)(x-a)+f(a) f '(x)=1/x, f '(a)=1 f(a)=ln(a)=0 So we get the linearization: y=1(x-1)+0=x-1